OpenStudy (anonymous):

2x + y = 3 x - 2y = -1 If equation one is multiplied by 2 and then the equations are added, the result is _____. 3x = 2 3x = 5 5x = 5

OpenStudy (anonymous):

4x+2y=6 +x-2y=-1 ------------ 5x=5